\begin{thebibliography}{11}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Bietti et~al.(2023)Bietti, Cabannes, Bouchacourt, Jegou, and Bottou]{bietti2023birth}
Alberto Bietti, Vivien Cabannes, Diane Bouchacourt, Herve Jegou, and Leon Bottou.
\newblock Birth of a transformer: A memory viewpoint.
\newblock In \emph{Advances in Neural Information Processing Systems 36}, 2023.

\bibitem[Brown et~al.(2020)Brown, Mann, Ryder, Subbiah, Kaplan, Dhariwal, Neelakantan, Shyam, Sastry, Askell, Agarwal, Herbert-Voss, Krueger, Henighan, Child, Ramesh, Ziegler, Wu, Winter, Hesse, Chen, Sigler, Litwin, Gray, Chess, Clark, Berner, McCandlish, Radford, Sutskever, and Amodei]{brown2020language}
Tom~B. Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel~M. Ziegler, Jeffrey Wu, Clemens Winter, Christopher Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei.
\newblock Language models are few-shot learners.
\newblock In \emph{Advances in Neural Information Processing Systems 33}, 2020.

\bibitem[Elhage et~al.(2021)Elhage, Nanda, Olsson, Henighan, Joseph, Mann, Askell, Bai, Chen, Conerly, DasSarma, Drain, Ganguli, Hatfield-Dodds, Hernandez, Jones, Kernion, Lovitt, Ndousse, Amodei, Brown, Clark, Kaplan, McCandlish, and Olah]{elhage2021mathematical}
Nelson Elhage, Neel Nanda, Catherine Olsson, Tom Henighan, Nicholas Joseph, Ben Mann, Amanda Askell, Yuntao Bai, Anna Chen, Tom Conerly, Nova DasSarma, Dawn Drain, Deep Ganguli, Zac Hatfield-Dodds, Danny Hernandez, Andy Jones, Jackson Kernion, Liane Lovitt, Kamal Ndousse, Dario Amodei, Tom Brown, Jack Clark, Jared Kaplan, Sam McCandlish, and Chris Olah.
\newblock A mathematical framework for transformer circuits.
\newblock \emph{Transformer Circuits Thread}, 2021.
\newblock https://transformer-circuits.pub/2021/framework/index.html.

\bibitem[Im et~al.(2023)Im, Kumar, Lattanzi, Moseley, and Vassilvitskii]{im2023massively}
Sungjin Im, Ravi Kumar, Silvio Lattanzi, Benjamin Moseley, and Sergei Vassilvitskii.
\newblock Massively parallel computation: Algorithms and applications.
\newblock \emph{Foundations and Trends{\textregistered} in Optimization}, 5\penalty0 (4):\penalty0 340--417, 2023.

\bibitem[Kushilevitz and Nisan(1997)]{kushilevitz1997communication}
Eyal Kushilevitz and Noam Nisan.
\newblock \emph{Communication Complexity}.
\newblock Cambridge University Press, 1997.

\bibitem[Olsson et~al.(2022)Olsson, Elhage, Nanda, Joseph, DasSarma, Henighan, Mann, Askell, Bai, Chen, Conerly, Drain, Ganguli, Hatfield-Dodds, Hernandez, Johnston, Jones, Kernion, Lovitt, Ndousse, Amodei, Brown, Clark, Kaplan, McCandlish, and Olah]{olsson2022context}
Catherine Olsson, Nelson Elhage, Neel Nanda, Nicholas Joseph, Nova DasSarma, Tom Henighan, Ben Mann, Amanda Askell, Yuntao Bai, Anna Chen, Tom Conerly, Dawn Drain, Deep Ganguli, Zac Hatfield-Dodds, Danny Hernandez, Scott Johnston, Andy Jones, Jackson Kernion, Liane Lovitt, Kamal Ndousse, Dario Amodei, Tom Brown, Jack Clark, Jared Kaplan, Sam McCandlish, and Chris Olah.
\newblock In-context learning and induction heads.
\newblock \emph{Transformer Circuits Thread}, 2022.
\newblock https://transformer-circuits.pub/2022/in-context-learning-and-induction-heads/index.html.

\bibitem[Peng et~al.(2024)Peng, Narayanan, and Papadimitriou]{peng2024limitations}
Binghui Peng, Srini Narayanan, and Christos Papadimitriou.
\newblock On limitations of the transformer architecture.
\newblock \emph{arXiv preprint arXiv:2402.08164}, 2024.

\bibitem[Radford et~al.(2019)Radford, Wu, Child, Luan, Amodei, and Sutskever]{radford2019language}
Alec Radford, Jeff Wu, Rewon Child, David Luan, Dario Amodei, and Ilya Sutskever.
\newblock Language models are unsupervised multitask learners.
\newblock Technical report, OpenAI, 2019.

\bibitem[Sanford et~al.(2023)Sanford, Hsu, and Telgarsky]{sanford2023representational}
Clayton Sanford, Daniel Hsu, and Matus Telgarsky.
\newblock Representational strengths and limitations of transformers.
\newblock In \emph{Advances in Neural Information Processing Systems 36}, 2023.

\bibitem[Sanford et~al.(2024)Sanford, Hsu, and Telgarsky]{sanford2024transformers}
Clayton Sanford, Daniel Hsu, and Matus Telgarsky.
\newblock Transformers, parallel computation, and logarithmic depth.
\newblock In \emph{Forty-First International Conference on Machine Learning}, 2024.

\bibitem[Vaswani et~al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin]{vaswani2017attention}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones, Aidan~N. Gomez, Lukasz Kaiser, and Illia Polosukhin.
\newblock Attention is all you need.
\newblock In \emph{Advances in Neural Information Processing Systems 30}, 2017.

\end{thebibliography}

\documentclass[11pt]{article}
\usepackage{fullpage}
\usepackage{amsmath,amsthm,amsfonts,amssymb,dsfont}

\usepackage[round]{natbib}

\usepackage{hyperref,cleveref}

\Crefformat{equation}{#2(#1)#3}
\crefformat{equation}{#2(#1)#3}
\Crefrangeformat{equation}{#3(#1)#4--#5(#2)#6}
\crefrangeformat{equation}{#3(#1)#4--#5(#2)#6}
\Crefmultiformat{equation}{#2(#1)#3}{ and~#2(#1)#3}{, #2(#1)#3}{ and~#2(#1)#3}
\crefmultiformat{equation}{#2(#1)#3}{ and~#2(#1)#3}{, #2(#1)#3}{ and~#2(#1)#3}

\usepackage{enumitem}
\setlist[description]{font=\normalfont}

\usepackage{csquotes}
\MakeOuterQuote{"}

\usepackage{mathtools}
\mathtoolsset{centercolon}
\DeclarePairedDelimiterXPP\ind[1]{\mathds{1}}{\lbrace}{\rbrace}{}{#1}
\DeclarePairedDelimiterX\eval[1]{\lbrace}{\rvert}{#1 \delimsize\rbrace}
\DeclarePairedDelimiter\ip{\langle}{\rangle}
\DeclarePairedDelimiter\abs{\lvert}{\rvert}
\DeclarePairedDelimiter\card{\lvert}{\rvert}

\DeclarePairedDelimiter\del{\lparen}{\rparen}

\DeclarePairedDelimiter\cbr{\lbrace}{\rbrace}
\DeclarePairedDelimiter\set{\lbrace}{\rbrace}

\DeclarePairedDelimiter\intcc{\lbrack}{\rbrack}

\newtheorem{theorem}{Theorem}

\DeclareMathOperator\softmax{softmax}

\newcommand\T{{\scriptscriptstyle{\mathsf{T}}}}
\newcommand\R{\mathbb{R}}
\newcommand\N{\mathbb{N}}

\newcommand\Sigmain{\Sigma_{\operatorname{in}}}
\newcommand\Sigmaout{\Sigma_{\operatorname{out}}}
\newcommand\encode{\psi_{\operatorname{in}}}
\newcommand\decode{\psi_{\operatorname{out}}}
\newcommand\TF{\mathsf{TF}}
\newcommand\SA{\mathsf{SA}}

\def\ddefloop#1{\ifx\ddefloop#1\else\ddef{#1}\expandafter\ddefloop\fi}
\def\ddef#1{\expandafter\def\csname bf#1\endcsname{\ensuremath{\mathbf{#1}}}}
\ddefloop abcdefghijklmnopqrstuvwxyzABCDEFGHIJKLMNOPQRSTUVWXYZ\ddefloop
\def\ddef#1{\expandafter\def\csname bf#1\endcsname{\ensuremath{\boldsymbol{\csname #1\endcsname}}}}
\ddefloop {alpha}{beta}{gamma}{delta}{epsilon}{varepsilon}{zeta}{eta}{theta}{vartheta}{iota}{kappa}{lambda}{mu}{nu}{xi}{pi}{varpi}{rho}{varrho}{sigma}{varsigma}{tau}{upsilon}{phi}{varphi}{chi}{psi}{omega}{Gamma}{Delta}{Theta}{Lambda}{Xi}{Pi}{Sigma}{varSigma}{Upsilon}{Phi}{Psi}{Omega}{ell}\ddefloop
\def\ddef#1{\expandafter\def\csname cal#1\endcsname{\ensuremath{\mathcal{#1}}}}
\ddefloop ABCDEFGHIJKLMNOPQRSTUVWXYZ\ddefloop

\title{One-layer transformers fail to solve the induction heads task}

\author{Clayton Sanford\thanks{Columbia University, \texttt{clayton@cs.columbia.edu}.} \and Daniel Hsu\thanks{Columbia University, \texttt{djhsu@cs.columbia.edu}.} \and Matus Telgarsky\thanks{New York University, \texttt{mjt10041@nyu.edu}.}}

\begin{document}
\maketitle

\begin{abstract}
  A simple communication complexity argument proves that no one-layer transformer can solve the induction heads task unless its size is exponentially larger than the size sufficient for a two-layer transformer.
\end{abstract}

\section{Introduction}

The mechanistic interpretability studies of \citet{elhage2021mathematical} and \citet{olsson2022context} identified the ubiquity and importance of so-called "induction heads" in transformer-based language models~\citep{vaswani2017attention,radford2019language,brown2020language}.
The basic task performed by an induction head is as follows.
\begin{itemize}
  \item The input is an $n$-tuple of tokens $(\sigma_1, \dotsc, \sigma_n)$ from a finite alphabet $\Sigma$.

  \item The output is an $n$-tuple of tokens $(\tau_1, \dotsc, \tau_n)$ from the augmented alphabet $\Sigma \cup \set{\bot}$, where the $i$-th output $\tau_i$ is equal to the input token that immediately follows the rightmost previous occurrence of the input token $\sigma_i$, or $\bot$ if there is no such previous occurrence.
    That is:
    $\tau_i = \bot$ if $\sigma_j \neq \sigma_i$ for all $j < i$, and otherwise
    $\tau_i = \sigma_{j_i+1}$ where $j_i = \max\set{ j : j < i \wedge \sigma_j = \sigma_i }$.
\end{itemize}
\citet{sanford2024transformers} call this the "$1$-hop induction heads task"; they also define and study generalizations of the task with increasing difficulty, which they call "$k$-hop" (for $k \in \N$).
The special case of $2$-hop is related to the function composition task defined and studied by \citet{peng2024limitations}.

\citet{bietti2023birth} gave an explicit construction of a transformer for solving the ($1$-hop) induction heads task.
Their construction is a two-layer transformer with a single attention head in each layer.
They also empirically found it difficult to train one-layer transformers to successfully solve the induction heads task under a certain data generation model, but training two-layer transformers was possible.
Indeed, \citet{elhage2021mathematical} noted: ``In small two-layer attention-only transformers, composition seems to be primarily used for one purpose: the creation of [\ldots] induction heads.''

In this note, we prove that a one-layer transformer cannot solve the induction heads task unless the size of the transformer is very large.
By "size", we mean the product of
the number of self-attention heads $h$,
the embedding dimension $m$, and
the number of bits of precision $p$ used by the transformer.
By "very large", we mean that $hmp = \Omega(n)$, where $n$ is the size of the input.
We note that when $\card{\Sigma} \leq n$, there is a two-layer transformer that solves the induction heads task with $h = O(1)$, $m = O(1)$, and $p = O(\log(n))$~\citep{bietti2023birth,sanford2024transformers}.
So our size lower bound for one-layer transformers is exponentially larger than the size that is sufficient for two-layer transformers.

The proof is based on a simple communication complexity argument.
Lower bounds on the size of one-layer transformers that solve related tasks were given by \citet{sanford2023representational} using similar arguments.
Conditional lower bounds for the $k$-hop (for general $k$, mentioned above) were given by \citet{sanford2024transformers} for $\Omega(\log k)$-layer transformers, assuming the 1-vs-2 cycle conjecture in the Massively Parallel Computation model~\citep{im2023massively}.
\citet{peng2024limitations} prove an average-case lower bound for one-layer transformers to solve their function composition task (which resembles the $2$-hop), again using a communication complexity argument.

\section{Transformer model}

In this section, we give a generic definition of one-layer transformers that allows for general forms of token embeddings and positional encodings.
A \emph{self-attention head} with \emph{embedding dimension} $m \in \N$ is a triple $(Q,K,V)$ where $Q \colon \N \times \R^m \to \R^m$, $K \colon \N \times \R^m \to \R^m$, and $V \colon \N \times \R^m \to \R^m$ are, respectively, called the query, key, and value embedding functions, and the first arguments to $Q, K, V$ enable the commonly-used positional encoding.
The self-attention head defines a mapping $\SA_{Q,K,V} \colon \R^{n \times m} \to \R^{n \times m}$ as follows.
On input $X = [ X_1, \dotsc, X_n ]^\T \in \R^{n \times m}$, the output $Y = [ Y_1, \dotsc, Y_n]^\T = \SA_{Q,K,V}(X) \in \R^{n \times m}$ is defined by
\begin{align*}
  Y_i
  & = \sum_{j=1}^n \alpha_{i,j} V(j,X_j) \\
  \quad \text{where} \quad
  (\alpha_{i,1},\dotsc,\alpha_{i,n})
  & = \softmax\del*{ \ip{Q(i,X_i), K(1,X_1)}, \dotsc, \ip{Q(i,X_i), K(n,X_n)} }
  \\
  & = \frac{\del*{ \exp\del{ \ip{Q(i,X_i), K(1,X_1)} }, \dotsc, \exp\del{ \ip{Q(i,X_i), K(n,X_n)} } }}{\sum_{j=1}^n \exp\del*{ \ip{Q(i,X_i), K(j,X_j)} }}
  .
\end{align*}
A \emph{one-layer transformer} with $h$ self-attention heads and embedding dimension $m$ is a collection of $h$ self-attention heads $(Q_t,K_t,V_t)_{t=1}^h$ each with embedding dimension $m$, together with an input encoding function $\encode \colon \N \times \Sigmain \to \R^m$ and an output decoding function $\decode \colon \N \times \R^m \to \Sigmaout$.
Here, the input alphabet $\Sigmain$ and the output alphabet $\Sigmaout$ are finite sets.
The transformer defines a mapping $\TF_{((Q_t,K_t,V_t)_{t=1}^h,\encode,\decode)} \colon \Sigmain^n \to \Sigmaout^n$ as follows.
On input $\sigma = (\sigma_1,\dotsc,\sigma_n) \in \Sigmain^n$, the output $\tau = (\tau_1,\dotsc,\tau_n) = \TF_{((Q_t,K_t,V_t)_{t=1}^h,\encode,\decode)}(\sigma) \in \Sigmaout^n$ is defined by
\begin{equation*}
  \tau_i = \decode\del*{ i, \encode\del*{ i, X_i} + Z_i }
  \quad \text{where} \quad
  [Z_1,\dotsc,Z_n]^\T = \sum_{t=1}^h \SA_{Q_t,K_t,V_t}\del*{ [\encode(1,X_1),\dotsc,\encode(n,X_n)]^\T } .
\end{equation*}
We say that a transformer uses \emph{$p$ bits of precision} if the outputs of all embedding functions ($Q_t$, $K_t$, $V_t$, $\encode$) and the self-attention heads may be rounded to rational numbers with at most $p$ bits of precision without changing the behavior of the mapping $\TF_{((Q_t,K_t,V_t)_{t=1}^h,\encode,\decode)}$.

\section{Size of one-layer transformers for the induction heads task}

\begin{theorem}
  If a one-layer transformer with $h$ self-attention heads, embedding dimension $m$, and $p$ bits of precision solves the induction heads task for input sequences of length $n$ over a three-symbol alphabet, then $hmp = \Omega(n)$.
\end{theorem}
\begin{proof}
  We give a reduction from the one-way communication problem INDEX~\citep[Example 4.19]{kushilevitz1997communication}.
  In this problem, Alice is given a bit string $f = (f_1,\dotsc,f_k) \in \{0,1\}^k$, and Bob is given an index $i^* \in [k]$.
  Alice can send a message to Bob, and after receiving it, Bob has to output $f_{i^*}$.
  By the pigeonhole principle, in any protocol for INDEX, Alice must send at least $k$ bits to Bob.

  Suppose there is a one-layer transformer (with $h$ self-attention heads and embedding dimension $m$, using $p$ bits of precision) that solves the induction heads task with a three-symbol alphabet $\Sigma = \set{0,1,?}$ (so $\Sigmain = \Sigma$ and $\Sigmaout = \Sigma \cup \set{\bot}$).
  We show that it specifies a one-way communication protocol for INDEX.
  Consider the input $n$-tuple $\sigma$, with $n=2k+1$, defined by
  \begin{equation*}
    \sigma = (e_1, f_1, e_2, f_2, \dotsc, e_k, f_k, ?) \in \set{0,1,?}^{2k+1}
  \end{equation*}
  where $f_1,\dotsc,f_k$ are taken from Alice's input, and
  \begin{equation*}
    e_i =
    \begin{cases}
      ? & \text{if $i = i^*$} , \\
      0 & \text{if $i \neq i^*$} ,
    \end{cases}
  \end{equation*}
  which is based on Bob's input.
  The $(2k+1)$-th output of the transformer on input $\sigma$ is $f_{i^*}$, which is exactly the correct output for INDEX.
  We show that Alice can send a message to Bob such that, subsequently, Bob can compute the $(2k+1)$-th output of the transformer and hence determine $f_{i^*}$.

  Consider one of the self-attention heads $(Q,K,V)$, and define $\tilde Q := Q \circ \encode$, $\tilde K = K \circ \encode$, and $\tilde V := V \circ \encode$.
  (We leave out the positional arguments in $Q, K, V, \encode$ for brevity.)
  The $(2k+1)$-th output of $\SA_{Q,K,V}((\encode(\sigma_1),\dotsc,\encode(\sigma_n)))$ is
  \begin{equation*}
    Y_{2k+1} =
    \frac{
      \underbrace{
        \sum_{i=1}^k \exp( \tilde Q(?)^\T \tilde K(e_i) ) \tilde V(e_i)
      }_{\text{known to Bob}}
      +
      \underbrace{
        \sum_{i=1}^k \exp( \tilde Q(?)^\T \tilde K(f_i) ) \tilde V(f_i)
      }_{\text{known to Alice}}
      +
      \underbrace{
        \exp( \tilde Q(?)^\T \tilde K(?) ) \tilde V(?)
      }_{\text{known to both}}
    }{
      \underbrace{
        \sum_{i=1}^k \exp( \tilde Q(?)^\T \tilde K(e_i) )
      }_{\text{known to Bob}}
      +
      \underbrace{
        \sum_{i=1}^k \exp(\tilde Q(?)^\T \tilde K(f_i) )
      }_{\text{known to Alice}}
      +
      \underbrace{
        \exp( \tilde Q(?)^\T \tilde K(?) )
      }_{\text{known to both}}
    }
    .
  \end{equation*}
  In order for Bob to compute $Y_{2k+1}$ to $p$ bits of precision,
  it suffices for Alice to send the values
  \begin{equation*}
    \frac
    {
      \sum_{i=1}^k \exp(\tilde Q(?)^\T \tilde K(f_i)) \tilde V(f_i)
    }
    {
      \sum_{i=1}^k \exp(\tilde Q(?)^\T \tilde K(f_i))
    }
    \in \R^m
    \quad\text{and}\quad
    \log\del*{ \sum_{i=1}^k \exp(\tilde Q(?)^\T \tilde K(f_i)) }
    \in \R
  \end{equation*}
  rounded to $O(p)$ bits of precision
  to Bob in a message of size $Cmp$ bits for some constant $C>0$ (see \Cref{sec:precision} for details).
  Such messages can be sent
  for all $h$ self-attention heads simultaneously; in total, Alice sends just $Chmp$ bits to Bob, and after that, Bob computes the output of the transformer and hence determines $f_{i^*}$, thereby solving the INDEX problem.
  Since every protocol for INDEX must require Alice to send at least $k$ bits, we have
  \begin{equation*}
    hmp \geq \frac{k}{C} = \frac{n-1}{2C} = \Omega(n)
    .
    \qedhere
  \end{equation*}
\end{proof}

\bibliographystyle{plainnat}
\bibliography{1hop}

\appendix

\section{Precision details}
\label{sec:precision}

The $j$-th component of $Y_{2k+1}$ can be expressed as $(A+B) / (Z_A + Z_B)$ where
\begin{align*}
  A
  & = \sum_{i=1}^k \exp( (\tilde Q(?)^\T \tilde K(f_i) ) \tilde V(f_i)_j
  ,
  & B
  & = \exp( \tilde Q(?)^\T \tilde K(?) ) \tilde V(?)_j + \sum_{i=1}^k \exp( \tilde Q(?)^\T \tilde K(e_i) ) \tilde V(e_i)_j
  ,
  \\
  Z_A
  & = \sum_{i=1}^k \exp( \tilde Q(?)^\T \tilde K(f_i) )
  ,
  & Z_B
  & = \exp( \tilde Q(?)^\T \tilde K(?) ) + \sum_{i=1}^k \exp( \tilde Q(?)^\T \tilde K(e_i) ) .
\end{align*}
Define $r := A / Z_A$ and $s := \log Z_A$.
Alice's message to Bob contains $\hat r$ and $\hat s$, which are obtained by rounding $r$ and $s$, respectively, to $3p$ bits of precision.
Hence, $\hat r$ and $\hat s$ satisfy $\abs{r - \hat r} \leq \epsilon$ and $\abs{s - \hat s} \leq \epsilon$ where $\epsilon := 2^{-3p}$.
It suffices to show that Bob can approximate $(A+B)/(Z_A+Z_B)$ up to error $2^{-p}$ using
\begin{equation*}
  \frac{\hat r e^{\hat s} + B}{e^{\hat s} + Z_B}
  ,
\end{equation*}
which only depends on $\hat r$, $\hat s$, $B$, and $Z_B$.
Observe that
\begin{equation*}
  \frac{\hat r e^{\hat s} + B}{e^{\hat s} + Z_B} - \frac{A+B}{Z_A+Z_B}
  = 
  \underbrace{
    r \del*{ \frac{e^{\hat s}}{e^{\hat s} + Z_B} - \frac{e^s}{e^s + Z_B} }
  }_{T_1}
  +
  \underbrace{
    \frac{\del{ \hat r - r } e^{\hat s}}{e^{\hat s} + Z_B}
  }_{T_2}
  +
  \underbrace{
    \frac{B}{Z_A + Z_B} \del*{ \frac{e^s - e^{\hat s}}{e^{\hat s} + Z_B} }
  }_{T_3}
  .
\end{equation*}
By assumption on the embedding functions, we may assume without loss of generality that $\abs{\tilde V(\sigma_i)_j} \leq 2^p$ for all $\sigma_i \in \Sigma$.
Therefore
\begin{equation*}
  \abs{r} \leq \max_{i\in[k]} \abs{\tilde V(f_i)_j} \leq 2^p
  \quad \text{and} \quad
  \frac{\abs{B}}{Z_A+Z_B} \leq \frac{\abs{B}}{Z_B} \leq \max\cbr*{ \abs{\tilde V(?)_j} , \max_{i\in[k]} \abs{\tilde V(e_i)_j} } \leq 2^p .
\end{equation*}
We now bound each of $T_1$, $T_2$, and $T_3$ in magnitude.
To bound $\abs{T_1}$, we use the $(1/4)$-Lipschitzness of the sigmoid function:
\begin{equation*}
  \abs{T_1}
  = \abs{r} \abs*{ \frac{e^{\hat s}}{e^{\hat s} + Z_B} - \frac{e^s}{e^s + Z_B} }
  \leq \frac{\abs{r} \abs{\hat s - s}}4
  \leq \frac{2^p\epsilon}{4} .
\end{equation*}
To bound $\abs{T_2}$:
\begin{equation*}
  \abs{T_2}
  = \abs{\hat r - r} \frac{e^{\hat s}}{e^{\hat s} + Z_B} \leq \abs{\hat r - r} \leq \epsilon .
\end{equation*}
Finally, to bound $\abs{T_3}$, we use the approximation $e^t (e^t - 1) \leq 1.25t$ for $t \in \intcc{0,1/8}$:
\begin{equation*}
  \abs{T_3}
  = \frac{\abs{B}}{Z_A + Z_B} \abs*{ \frac{e^s - e^{\hat s}}{e^{\hat s} + Z_B} }
  \leq \frac{\abs{B}}{Z_A + Z_B} \cdot \frac{Z_A}{Z_A+Z_B} \cdot \frac{e^{\abs{\hat s - s}} - 1}{e^{-\abs{\hat s - s}}}
  \leq 2^p \cdot 1.25\epsilon .
\end{equation*}
Therefore
\begin{equation*}
  \abs*{
    \frac{\hat r e^{\hat s} + B}{e^{\hat s} + Z_B}  - \frac{A+B}{Z_A+Z_B}
  }
  \leq \abs{T_1} + \abs{T_2} + \abs{T_3}
  \leq \del*{ \frac32 \cdot 2^p + 1 } \epsilon
  \leq 2^{-p}
  .
\end{equation*}

\end{document}